OpenStudy (anonymous):

consider the curve y^2 = 2 + (xy) find all points on the curve where the line tangent to the curve has a slope of 1/2

OpenStudy (anonymous):

\[y ^{2} = 2 + (xy)\]

hero (hero):

I'll try to get the solution to this for you.

hero (hero):

Actually, mimi should know. Help him mimi

OpenStudy (mimi_x3):

hm, i don't know sorry. i thought i knew :/

hero (hero):

So mimi, you don't know basic math or calculus

OpenStudy (anonymous):

:o this one has my entire class stumped and the teacher wont give us the answer

OpenStudy (mimi_x3):

this looks like implicit differentiation. I know calculus but not this.

OpenStudy (anonymous):

yeah you have to implicitly differentiate the equation then set it equal to .5 i think but not sure where to go after that

hero (hero):

If you know how to get to that point, then do it. Maybe we can help you from there.

OpenStudy (anonymous):

\[dy/dx = y /(2y-x)\] thing is they dont give you anything to find a point on the graph so im thinking if you solve the original for x and substitute it might work

hero (hero):

That sounds like a good idea

OpenStudy (anonymous):

\[1/2 = y /(2y - (y ^{2} -2 /y))\]

OpenStudy (anonymous):

i get stuck about here because I cant algebra and / or you cant do that XD

hero (hero):

I see, well, hopefully, what you've done is correct so far.

hero (hero):

After doing the math I get \[y = \sqrt[3]{2}\] I can show you what I did

OpenStudy (anonymous):

ok

hero (hero):

I will post a vyew. You're going to laugh when you see how easy it is

OpenStudy (anonymous):

haha it probably is pretty easy i just get hung up on basic math like when you have like x + 3/x^2

hero (hero):

You did a good job too. Couldn't have done it without you :P

OpenStudy (anonymous):

The coordinates of the solution points are:\[\left(0,\sqrt{2}\right),\left(0,-\sqrt{2}\right) \]Refer to the Mathematica attachment.